OpenStudy (anonymous):

Please help me solve n +5 3/10 = 8 1/10. I am in 4th grade. I need help trying to find the value of n. Thanks!

OpenStudy (anonymous):

n=8 1/10 - 5 3/10 n=81/10 - 53/10 n=28/10 n=2.8